OpenStudy (kirbz):

What is the value of 2x to the second power + 18 when x = -6

OpenStudy (xapproachesinfinity):

is \((2x)^{2}+18\)

OpenStudy (xapproachesinfinity):

is it like that?

OpenStudy (wendy.ivette11714):

\[2(−6)^2+18\] \[= ?\] @Kirbz

OpenStudy (kirbz):

xapproache. and x = negative 6

OpenStudy (xapproachesinfinity):

or is it \(2x^2+18\)

OpenStudy (anonymous):

wendy is on the right track

OpenStudy (kirbz):

the second one xapproache

OpenStudy (xapproachesinfinity):

yes but i want to make sure what you mean by 2x to the power 2 +18

OpenStudy (xapproachesinfinity):

ok then replace x by -6

OpenStudy (anonymous):

the equation should read 2(-6)^2+18

OpenStudy (xapproachesinfinity):

#themazinmathmaster did it already just calculate

OpenStudy (anonymous):

then simplify

OpenStudy (kirbz):

thanks guys!

OpenStudy (xapproachesinfinity):

(-6)^2=?

OpenStudy (anonymous):

-6*6

OpenStudy (anonymous):

-6*-6 my bad

OpenStudy (anonymous):

what does that equal kirby

OpenStudy (kirbz):

36

OpenStudy (anonymous):

ok so now u have 2*36+18=?

OpenStudy (kirbz):

88?

OpenStudy (kirbz):

90.

OpenStudy (kirbz):

sorry

OpenStudy (anonymous):

yyes

OpenStudy (anonymous):

medal?

OpenStudy (kirbz):

:D